(5 pts)  Suppose I sell 10 CALL option contracts on Coke with a strike price of $47.00...

  1. (5 pts)  Suppose I sell 10 CALL option contracts on Coke with a strike price of $47.00 and an option premium of 26 cents. The stock price is $45.83 when I sell the call and moves between $43 and $46.25 during the time up to expiration. Assuming that I do not own any Coke stock, what is my profit or loss on this?
  1. (5 pts) Using the date in question 11, what would my profit or loss be if the price of Coke stock moved from $40 and $49.75?
Answer #1

1.
Profit at 43=-10*100*(MAX(43-47,0)-0.26)=260
Profit at 47=-10*100*(MAX(46.25-47,0)-0.26)=9500

Hence, maximum profit is 260

2.
Profit at 40=-10*100*(MAX(40-47,0)-0.26)=260
Profit at 49.75=-10*100*(MAX(49.75-47,0)-0.26)=-2490

Hence, maximum profit is 260
maximum loss is 2490